\begin{thebibliography}{18}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Agi{\'c} \& Vuli{\'c}(2019)Agi{\'c} and
  Vuli{\'c}]{agic-vulic-2019-jw300}
{\v{Z}}eljko Agi{\'c} and Ivan Vuli{\'c}.
\newblock {JW}300: A wide-coverage parallel corpus for low-resource languages.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pp.\  3204--3210, Florence, Italy, July 2019.

\bibitem[Bahdanau et~al.(2015)Bahdanau, Cho, and
  Bengio]{DBLP:journals/corr/BahdanauCB14}
Dzmitry Bahdanau, Kyunghyun Cho, and Yoshua Bengio.
\newblock Neural machine translation by jointly learning to align and
  translate.
\newblock In Yoshua Bengio and Yann LeCun (eds.), \emph{3rd International
  Conference on Learning Representations, {ICLR} 2015, San Diego, CA, USA, May
  7-9, 2015, Conference Track Proceedings}, 2015.

\bibitem[Christodouloupoulos \& Steedman(2015)Christodouloupoulos and
  Steedman]{Christodouloupoulos2015}
Christos Christodouloupoulos and Mark Steedman.
\newblock A massively parallel corpus: the bible in 100 languages.
\newblock \emph{Language Resources and Evaluation}, 49\penalty0 (2):\penalty0
  375--395, Jun 2015.

\bibitem[Dong et~al.(2015)Dong, Wu, He, Yu, and Wang]{dong-etal-2015-multi}
Daxiang Dong, Hua Wu, Wei He, Dianhai Yu, and Haifeng Wang.
\newblock Multi-task learning for multiple language translation.
\newblock In \emph{Proceedings of the 53rd Annual Meeting of the Association
  for Computational Linguistics and the 7th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pp.\  1723--1732,
  Beijing, China, july 2015.

\bibitem[Kalchbrenner \& Blunsom(2013)Kalchbrenner and
  Blunsom]{kalchbrenner-blunsom-2013-recurrent-continuous}
Nal Kalchbrenner and Phil Blunsom.
\newblock Recurrent continuous translation models.
\newblock In \emph{Proceedings of the 2013 Conference on Empirical Methods in
  Natural Language Processing}, pp.\  1700--1709, Seattle, Washington, USA,
  October 2013. Association for Computational Linguistics.

\bibitem[Kingma \& Ba(2014)Kingma and Ba]{Kingma2014AdamAM}
Diederik~P. Kingma and Jimmy Ba.
\newblock Adam: A method for stochastic optimization.
\newblock \emph{CoRR}, abs/1412.6980, 2014.

\bibitem[Klein et~al.(2018)Klein, Kim, Deng, Nguyen, Senellart, and
  Rush]{klein-etal-2018-opennmt}
Guillaume Klein, Yoon Kim, Yuntian Deng, Vincent Nguyen, Jean Senellart, and
  Alexander Rush.
\newblock {O}pen{NMT}: Neural machine translation toolkit.
\newblock In \emph{Proceedings of the 13th Conference of the Association for
  Machine Translation in the {A}mericas (Volume 1: Research Papers)}, pp.\
  177--184, Boston, MA, March 2018.

\bibitem[Koehn et~al.(2007)Koehn, Hoang, Birch, Callison-Burch, Federico,
  Bertoldi, Cowan, Shen, Moran, Zens, Dyer, Bojar, Constantin, and
  Herbst]{koehn-etal-2007-moses}
Philipp Koehn, Hieu Hoang, Alexandra Birch, Chris Callison-Burch, Marcello
  Federico, Nicola Bertoldi, Brooke Cowan, Wade Shen, Christine Moran, Richard
  Zens, Chris Dyer, Ond{\v{r}}ej Bojar, Alexandra Constantin, and Evan Herbst.
\newblock {M}oses: Open source toolkit for statistical machine translation.
\newblock In \emph{Proceedings of the 45th Annual Meeting of the Association
  for Computational Linguistics Companion Volume Proceedings of the Demo and
  Poster Sessions}, pp.\  177--180, Prague, Czech Republic, June 2007.

\bibitem[Lavie \& Agarwal(2007)Lavie and Agarwal]{meteor}
Alon Lavie and Abhaya Agarwal.
\newblock Meteor: An automatic metric for mt evaluation with high levels of
  correlation with human judgments.
\newblock In \emph{Proceedings of the Second Workshop on Statistical Machine
  Translation}, StatMT ’07, pp.\  228–231, USA, 2007.

\bibitem[Neubig \& Hu(2018)Neubig and Hu]{neubig-hu-2018-rapid}
Graham Neubig and Junjie Hu.
\newblock Rapid adaptation of neural machine translation to new languages.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing}, pp.\  875--880, Brussels, Belgium,
  October-November 2018.

\bibitem[Papineni et~al.(2002)Papineni, Roukos, Ward, and
  Zhu]{papineni-etal-2002-bleu}
Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu.
\newblock {B}leu: a method for automatic evaluation of machine translation.
\newblock In \emph{Proceedings of the 40th Annual Meeting of the Association
  for Computational Linguistics}, pp.\  311--318, Philadelphia, Pennsylvania,
  USA, July 2002.

\bibitem[Popovi{\'c}(2015)]{popovic-2015-chrf}
Maja Popovi{\'c}.
\newblock chr{F}: Character n-gram {F}-score for automatic {MT} evaluation.
\newblock In \emph{Proceedings of the Tenth Workshop on Statistical Machine
  Translation}, pp.\  392--395, Lisbon, Portugal, September 2015.

\bibitem[Sennrich et~al.(2016)Sennrich, Haddow, and
  Birch]{sennrich-etal-2016-neural}
Rico Sennrich, Barry Haddow, and Alexandra Birch.
\newblock Neural machine translation of rare words with subword units.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pp.\  1715--1725,
  Berlin, Germany, August 2016.

\bibitem[Tedla \& Yamamoto(2016)Tedla and Yamamoto]{Tedla2016TheEO}
Yemane Tedla and Kazuhide Yamamoto.
\newblock The effect of shallow segmentation on english-tigrinya statistical
  machine translation.
\newblock \emph{2016 International Conference on Asian Language Processing
  (IALP)}, pp.\  79--82, 2016.

\bibitem[Teferra~Abate et~al.(2018)Teferra~Abate, Melese, Yifiru~Tachbelie,
  Meshesha, Atinafu, Mulugeta, Assabie, Abera, Ephrem, Abebe, Tsegaye, Lemma,
  Andargie, and Shifaw]{teferra-abate-etal-2018-parallel}
Solomon Teferra~Abate, Michael Melese, Martha Yifiru~Tachbelie, Million
  Meshesha, Solomon Atinafu, Wondwossen Mulugeta, Yaregal Assabie, Hafte Abera,
  Binyam Ephrem, Tewodros Abebe, Wondimagegnhue Tsegaye, Amanuel Lemma, Tsegaye
  Andargie, and Seifedin Shifaw.
\newblock Parallel corpora for bi-directional statistical machine translation
  for seven {E}thiopian language pairs.
\newblock In \emph{Proceedings of the First Workshop on Linguistic Resources
  for Natural Language Processing}, pp.\  83--90, Santa Fe, New Mexico, USA,
  August 2018.

\bibitem[Tiedemann(2012)]{TIEDEMANN12.463}
Jörg Tiedemann.
\newblock Parallel data, tools and interfaces in opus.
\newblock In \emph{Proceedings of the Eight International Conference on
  Language Resources and Evaluation (LREC'12)}, Istanbul, Turkey, May 2012.

\bibitem[Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez,
  Kaiser, and Polosukhin]{DBLP:conf/nips/VaswaniSPUJGKP17}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N. Gomez, Lukasz Kaiser, and Illia Polosukhin.
\newblock Attention is all you need.
\newblock In Isabelle Guyon, Ulrike von Luxburg, Samy Bengio, Hanna~M. Wallach,
  Rob Fergus, S.~V.~N. Vishwanathan, and Roman Garnett (eds.), \emph{Advances
  in Neural Information Processing Systems 30: Annual Conference on Neural
  Information Processing Systems 2017, 4-9 December 2017, Long Beach, CA,
  {USA}}, pp.\  5998--6008, 2017.

\bibitem[Zoph et~al.(2016)Zoph, Yuret, May, and
  Knight]{zoph-etal-2016-transfer}
Barret Zoph, Deniz Yuret, Jonathan May, and Kevin Knight.
\newblock Transfer learning for low-resource neural machine translation.
\newblock In \emph{Proceedings of the 2016 Conference on Empirical Methods in
  Natural Language Processing}, pp.\  1568--1575, Austin, Texas, November 2016.

\end{thebibliography}

\documentclass{article} 
\usepackage{iclr2020_conference,times}

\usepackage{amsmath,amsfonts,bm}

\def\eqref#1{equation~\ref{#1}}

\def\1{\bm{1}}

\DeclareMathAlphabet{\mathsfit}{\encodingdefault}{\sfdefault}{m}{sl}
\SetMathAlphabet{\mathsfit}{bold}{\encodingdefault}{\sfdefault}{bx}{n}

\usepackage{hyperref}
\usepackage{url}

\title{Tigrinya Neural Machine Translation with Transfer Learning for Humanitarian Response}

\author{Alp \"Oktem, Mirko Plitt \& Grace Tang \\
Translators without Borders\\
30 Main Street STE 500\\
Danbury, CT 06810, USA \\
\texttt{\{alp,mirko,grace\}@translatorswithoutborders.org} \\
}

\iclrfinalcopy 
\begin{document}

\maketitle
\setcounter{footnote}{-1}
\begin{abstract}
We report our experiments in building a domain-specific Tigrinya-to-English neural machine translation system. We use transfer learning from other Ge'ez script languages and report an improvement of 1.3 BLEU points over a classic neural baseline. We publish our development pipeline as an open-source library and also provide a demonstration application.
\end{abstract}

\section{Introduction}

Tigrinya (also spelled Tigrigna) is an Ethiopic language spoken by around 7.9 million people in Eritrea and Ethiopia. It is neither supported by any commercial machine translation (MT) provider, nor has any publicly available models. Refugees who speak Tigrinya face language and communication barriers when arriving in Europe. An MT system could improve access to  information and enable two-way communication so refugees have a voice and can share their needs.

The complex morphological structure of Tigrinya makes it especially challenging for statistical MT \citep{Tedla2016TheEO, teferra-abate-etal-2018-parallel}. Neural MT, on the other hand, can overcome these problems with methods like subword segmentation \citep{sennrich-etal-2016-neural} and lead to more accurate models \citep{kalchbrenner-blunsom-2013-recurrent-continuous, DBLP:journals/corr/BahdanauCB14}. Known to be a data-hungry technology, it is now possible to train neural-based MT for Tigrinya thanks to recently released public datasets \citep{agic-vulic-2019-jw300, teferra-abate-etal-2018-parallel}. An advantage of neural MT is the availability of techniques like cross-lingual transfer learning \citep{zoph-etal-2016-transfer} and multilingual training \citep{dong-etal-2015-multi} which help leverage data from other languages and are especially suitable in low-resource scenarios \citep{neubig-hu-2018-rapid}.

In this paper, we explain the development of a Tigrinya-to-English neural MT model using publicly available datasets in Ge'ez-scripted languages. Our models are further adapted to the humanitarian domain to improve the translation capabilities of Translators without Borders (TWB), a non-profit organization offering language and translation support for humanitarian and development agencies, and other non-profit organizations.

\vspace*{-1mm}

\section{Experiments}
\vspace*{-1mm}
\subsection{Data}

We gathered an internal dataset from sentences in TWB's translation memories. This dataset is both used for in-domain training and for testing. Two hundred sentences of varying lengths were selected randomly as a test set. This and other sources of public parallel corpora used in this work are listed in Table \ref{data-table}.

\begin{table}[t]
\caption[caption]{Parallel corpora used in this work. Dataset names marked with an asterisk are available through OPUS repository \citep{TIEDEMANN12.463, Christodouloupoulos2015}. Ethiopian languages corpus \citep{teferra-abate-etal-2018-parallel} is also openly available online\footnotemark.}
\label{data-table}
\begin{center}
\begin{tabular}{l|cccccccc}
  & JW300* & Ethiopian & Bible- & Global & GNOME* & Tanzil* & TWB & \textbf{TOTAL}\\
  &  & corpus & uedin* & Voices* &  &  & 
\\ \hline \\
Amharic  & 722K  & 66K               & 61K         & 1,6K          & 57K   & 94K    & - & 1M   \\
Ge'ez    & -     & 11K               & -           & -             & -     & -      & -   & 11K  \\
Tigrinya & 400K  & 36K               & -           & -             & -     & -      & 2.5K & 439K \\
\end{tabular}
\end{center}
\end{table}

\vspace*{-2mm}
\subsection{Experimental setup}

The transfer-learning-based training process consists of three stages. First, we train the model on a shuffled mix of all datasets totaling up to 1.45 million sentences. Second, we fine-tune the model on Tigrinya using only the Tigrinya portion of the mix (438,000 sentences). In the third phase, we fine-tune on the training partition of our in-house data (2,300 sentences). As a baseline, we skip the first multilingual training step and use only Tigrinya data. The model is later fine-tuned to in-domain data in the same way.

OpenNMT-py toolkit \citep{klein-etal-2018-opennmt} is used for training the models. The model consists of an 8-head Transformer \citep{DBLP:conf/nips/VaswaniSPUJGKP17} with 6-layer hidden units of 512 unit size. A token-batch size of 4096, 2048 and 10 was selected for multilingual, unilingual and in-domain training respectively. As for the optimizer, Adam \citep{Kingma2014AdamAM} was chosen with 4000 warm-up steps. Trainings were performed until no further improvement was recorded in development set perplexity in the last 5 validations. This resulted in 73,500, 85,000 and 85,240 steps for each stage.

\footnotetext{\url{http://github.com/AAUThematic4LT/Parallel-Corpora-for-Ethiopian-Languages}}

Byte-pair encoding (BPE) models were trained separately for Latin script and Ge'ez script using 6,000 steps. English sentences were lowercased and tokenized beforehand using Moses tokenizer \citep{koehn-etal-2007-moses}. Ge'ez-scripted sentences were tokenized using a punctuation separation script\footnote{\url{http://github.com/translatorswb/mt-tools}}.

\vspace*{-2mm}

\subsection{Results}

We report our test set scores at each stage together with the baseline using various commonly used automatic evaluation metrics in Table \ref{results}. Results show an agreement between all evaluation measures on the boost obtained from multilingual pre-training. Accuracy increases of +1.3, +3.1 and +0.9 points are recorded using \textit{BLEU} \citep{papineni-etal-2002-bleu}, \textit{ChrF} \citep{popovic-2015-chrf} and \textit{Meteor} \citep{meteor} metrics respectively.
\vspace*{-2mm}
\begin{table}[t]
\caption{Automatic evaluation results for baseline approach and at each stages of our training pipeline.}.
\label{results}
\begin{center}
\begin{tabular}{l|ccc}
  & BLEU  & ChrF  & Meteor  \\ \hline \\
Baseline  & 22.28 &	46.51 &	26.1 \\
Multilingual   & 15.84 &	40.99 &	23.32  \\
Tigrinya  & 17.8 &	42.92 &	24.61 \\
In-domain & \textbf{23.6} &	\textbf{49.59} & \textbf{27.04} \\
\end{tabular}
\end{center}
\end{table}

\vspace*{-2mm}

\section{Conclusion}
\vspace*{-1mm}

With this work, we have demonstrated the utility of cross-lingual transfer learning on building a Tigrinya-to-English MT system. As a result of this work, a demonstration application was launched as the first neural Tigrinya-to-English translator\footnote{\url{http://gamayun.translatorswb.org/tigrinya}}. As for future work, we will develop English-to-Tigrinya models and evaluate the usability of the bidirectional system in a humanitarian setting using feedback from native speakers.

\subsubsection*{Acknowledgments}
This work was done partially in collaboration with the Masakhane initiative\footnote{\url{http://www.masakhane.io/}}. Special thanks to Musie Meressa Berhe for helping revise our dataset.

\bibliography{iclr2020_conference}
\bibliographystyle{iclr2020_conference}

\end{document}